Science Report Apr 14th

Science Report 14-Apr-2021

· Extremophiles Research

o Collected 2 samples while on EVA, and samples stored in fridge in the Science Dome. Intended to collect 4 samples, so at a later date, a following EVA might need to return to the coordinates of this collection site.

· Hydroponic Garden

o Significant leaking from the compost and H2O (control) pipes. Monitoring with catch buckets underneath. Still unable to measure evaporation due to leak.

o Began to collect data on stalk height and root growth

o Green Hab heater to be turned on at approx. 5pm every evening with temperature set to 65F.

In addition, after this report is sent, Crew Botanist will be planting seeds this evening.

Olivia Ettlin & Julio Hernandez

Copyright © The Mars Society. All rights reserved. | Main Site